OpenStudy (anonymous):

Consider the functions f(x) = 3x − 2 and g(x) = x2 − 1. What is the value of f[g(5)]? 312 70 168 −1

OpenStudy (amorfide):

okay so f[g(5)] this means you start with the function g so you do g(5) which you know is substituting x=5 into the function g once you do this you will have f(g) g is the value you get from the function g after substituting x=5 you would substitute that value into f

OpenStudy (anonymous):

so f+g 3(5)-2=13 (5)2-1=9

OpenStudy (anonymous):

13-9=4

OpenStudy (anonymous):

???

OpenStudy (anonymous):

When you wrote "x2 - 1" for g(x), did you switch the order of x and 2 or is g(x) = x^2 + 1? Is x squared?

OpenStudy (anonymous):

squared

OpenStudy (anonymous):

3(5)-2=13 (5)2-1=24

OpenStudy (anonymous):

OK. Then as amorfide said, first substitute 5 for x in x^2 - 1 to get the value of g(x). But the you are supposed to to use g(5) = 24 for the value passed to f(x), in f(g(5)) the x will be equal to g(5).

OpenStudy (anonymous):

??

OpenStudy (anonymous):

the x in each function represents the value passed to the equation. x is not the same value for each. For g(x) x = 5, but for f(g(x)) the x in f(x), x = g(5).